# Self-supervised visual learning in the low-data regime: a comparative evaluation

### v1

@@ 5 Conclusions
In this paper, the issue of Self-Supervised Learning (SSL) in real-world application settings, i.e., cases where it is not always feasible to collect and/or to utilize very large (in the order of millions of instances) pretraining datasets, was examined. In this context, this work introduced a taxonomy of modern SSL methods for image analysis, accompanied by detailed explanations and insights regarding the main categories of approaches. Subsequently, a comprehensive comparative experimental study was performed in the so-called ‘low-data regime’, i.e., assuming a pretraining dataset with  50k -  300k images, aiming at shedding light on how the various SSL methods behave in such training scenarios. The performed comparative evaluation showcased that a pretraining dataset size limit exists, in the range  50k -  65k images, below which SSL is ineffective for representation learning. In the majority of cases, supervised pretraining and large-scale DINOv2 pretraining surpassed low-data SSL pretraining in terms of both downstream accuracy and robustness, with DINO retaining its ability to learn more generally applicable embeddings in the low-data regime. Very interestingly, in domain-specific experiments, in-domain low-data SSL pretraining outperformed all competitors and even large-scale pretraining on general datasets. This finding suggests that when extreme amounts of data are not available and/or the target downstream dataset is significantly different in nature compared to the large-scale datasets commonly used for pretraining (e.g., medical and security imaging against conventional natural RGB images), in-domain low-data SSL pretraining offers recognition advantages. At a finer level of detail, when comparing the performance of SSL methods only, MAE seems to need less pretraining data than DINO, in order to learn generally applicable features that are inherently semantically discriminant and consistent with the visible objects. On the other hand, given a reasonably-sized downstream dataset for finetuning, DINO typically surpasses MAE in the low-data setting, since it achieves higher downstream classification accuracy and generates the most well-separated and compact class representations. Moreover, the superiority of DINO persisted in domain-specific scenarios, when moving beyond the RGB space, where it outperformed all SSL competitors and even large-scale supervised pretraining. This indicates that low-data SSL pretraining not only works, but is also highly relevant in cases of domain-specific downstream tasks where generic (e.g., ImageNet-1k pretraining-derived) representations cannot generalize well.

### v2

@@ 5 Conclusions
In this paper, the issue of Self-Supervised Learning (SSL) in real-world application settings, i.e., cases where it is not always feasible to collect and/or to utilize very large (in the order of millions of instances) pretraining datasets, was examined. In this context, based on a taxonomy of modern SSL methods for image analysis, a comprehensive comparative experimental study was performed in the so-called ‘low-data regime’, i.e., assuming a pretraining dataset with  50k -  300k images, aiming at shedding light on how the various SSL methods behave in such training scenarios. The performed comparative evaluation showcased that a pretraining dataset size limit exists, in the range  50k -  65k images (assuming identical pretraining and downstream image domains), below which SSL is ineffective for representation learning. In the majority of cases, supervised pretraining and large-scale DINOv2 pretraining surpassed low-data SSL pretraining in terms of both downstream accuracy and robustness, with DINO retaining its ability to learn more generally applicable embeddings in the low-data regime. Very interestingly, in domain-specific experiments, in-domain low-data SSL pretraining outperformed all competitors and even large-scale pretraining on general datasets. This finding suggests that when extreme amounts of data are not available and/or the target downstream dataset is significantly different in nature compared to the large-scale datasets commonly used for pretraining (e.g., medical and security imaging against conventional natural RGB images), in-domain low-data SSL pretraining offers recognition advantages. At a finer level of detail, when comparing the performance of SSL methods only, MAE seems to need less pretraining data than DINO, in order to learn generally applicable features that are inherently semantically discriminant and consistent with the visible objects. On the other hand, given a reasonably-sized downstream dataset for finetuning, DINO typically surpasses MAE in the low-data setting, since it achieves higher downstream classification accuracy and generates the most well-separated and compact class representations. Moreover, the superiority of DINO persisted in domain-specific scenarios, when moving beyond the RGB space, where it outperformed all SSL competitors and even large-scale supervised pretraining. This indicates that low-data SSL pretraining not only works, but is also highly relevant in cases of domain-specific downstream tasks where generic (e.g., ImageNet-1k pretraining-derived) representations cannot generalize well.
@@ Acknowledgment
Appendix A Details of SSL pretext tasks
This Appendix details a selection of important pretext tasks for each of the 4 main visual SSL categories: generative, contrastive, self-distillation and clustering approaches.
A.1 Generative pretext
A.1.1 Rotation Prediction
$\mathcal{L}_{Rot}=\mathcal{L}_{CE}(\mathbf{y},\mathbf{p}),$ (1)
A.1.2 Colorization
$\mathcal{L}_{Col}=\mathcal{L}_{MSE}(\mathbf{P},\mathbf{Y}).$ (2)
A.1.3 Super-resolution
$\mathcal{L}_{SR}=\mathcal{L}_{MSE}(\mathbf{P},\mathbf{Y}).$ (3)
A.1.4 BEiT
The loss function employed for BEiT pretraining is:
6:  Compute $\mathcal{L}_{BEiT}$.
A.1.5 MAE
$\mathcal{L}_{Mas}=\mathcal{L}_{MSE}(\mathbf{P},\mathbf{Y}).$ (5)
4:  Compute loss $\mathcal{L}_{Mas}(\mathbf{P},\mathbf{Y})$.
A.1.6 I-JEPA
[…]
A.2 Contrastive pretext
A.2.1 Early methods
A.2.2 MoCo
A.2.3 PIRL
A.2.4 SimCLR
Algorithm 3 SimCLR pseudocode for a single mini-batch
1:  for each $\mathbf{B}^{i}\in\mathcal{B}$ do
5:  end for
6:  for each $i=1$ to $B$ do
7:     Compute $\mathcal{L}_{NT-Xent}(\mathcal{B})$.
9:  end for7
A.2.5 Barlow Twins
A.3 Self-distillation pretext
A.3.1 BYOL
$\mathcal{L}_{BYOL}=\mathcal{L}_{MSE}(\mathbf{b}_{1},\mathbf{b}_{2}).$ (10)
A.3.2 SimSiam
A.3.3 DINO
Algorithm 4 DINO pretext task pseudocode for a single training image
5:  Compute $\mathcal{L}_{DINO}$.
  • 1.
  • 2.
  • 3.
A.4 Clustering pretext
A.4.1 DeepCluster
Algorithm 5 DeepCluster pretext task pseudocode
1:  Initialize parameters of $f(\cdot)$ (e.g., randomly).
2:  while not converged do
5:     for each mini-batch $\mathcal{B}$ sampled from $\mathcal{X}$ do
6:        Compute $\mathcal{L}_{DeepCluster}$ for each image in $\mathcal{B}$.
8:     end for
9:  end while
  • 1.
  • 2.
  • 3.
  • 4.
A.4.2 SeLa
s.t.:
$\sum_{y=1}^{K}q(y|\mathbf{X}_{i})=1,\forall i\in{1,2,\dots,N},$ (16)
A.4.3 SwAV
[…]
$\mathbf{p}^{v*}_{i}=softmax(\mathbf{C}\mathbf{h}^{v*}_{i}),$ (19)